# TNO, The Hague, The Netherlands

### v1

@@ Abstract
The paper provides an overview of quantum approaches to routing, logistic network design, fleet maintenance, cargo loading, prediction, and scheduling problems. Notably, most solutions in the literature are hybrid, combining quantum and classical computing. The conclusion highlights the early stage of quantum computing, emphasising its potential impact on logistics and supply chain optimisation. In the final overview, the literature is categorised, identifying quantum annealing dominance and a need for more research in prediction and machine learning is highlighted. The consensus is that quantum computing has great potential but faces current hardware limitations, necessitating further advancements for practical implementation.
@@
Quantum Computing Logistcs Supply Chain Management Optimisation
@@ 1 Introduction
Quantitative optimisation is integral to logistics and supply chain management, driving efficiency, cost reduction, and overall performance improvement. Various challenges underscore the need for this field. These kind of challenges and the corresponding optimisation problems are often divided into multiple levels, like operational, tactical and strategic brouer2018optimisation ; de2017framework , see Fig.  1. On an operational level, there is efficient route planning, as seen in vehicle routing problems (VRP), minimising transportation costs while meeting demand and adhering to constraints. Also problems like stowage and cargo loading have to be solved on an operational level. Next, in inventory management, a balance is sought between over-stocking and under-stocking, employing models like Economic Order Quantity (EOQ) for optimal order quantities. Production schedule optimisation addresses job scheduling, employee and machine scheduling, and production planning, minimising lead times and costs. Last, accurate demand forecasting is needed to optimise inventory, production, and transportation planning. On a tactical level, we have problems like fleet deployment, timetabling and hub site design. On a strategic level, decisions on facility location, fleet size and mix, and network design, impact logistics costs.
Next to these pure optimisation problems, there are two issues we also have to address in quantitative optimisation in this field. First, supply chain objectives often conflict, requiring multi-objective optimisation for informed decisions. Supplier selection involves balancing factors like cost, quality, lead times, and reliability. Risk management minimises disruptions from natural disasters or political events, ensuring operational continuity. Second, in logistics we have to cope with the dynamic behaviour of the system. Dynamic optimisation adapts to real-time changes, crucial in a dynamic supply chain environment. Green supply chain optimisation focuses on sustainability, reducing environmental impact through carbon emission and energy usage minimisation.
@@
These issues make decision-making processes even more complex and consider various variables and constraints, ensuring a smooth flow of goods while maximising operational efficiency and minimising costs. Advanced techniques, including mathematical modelling and metaheuristics, are commonly used to tackle these challenges, which are computationally hard in most cases.
@@
Currently, there is a burgeoning computing paradigm known as quantum computing, which holds the potential to transform the contemporary computing landscape. Quantum computers, the devices central to this paradigm, leverage principles from quantum mechanics to achieve computational speeds and efficiencies that surpass the inherent capabilities of current classical computers for specific problems. The notable advancements attributable to quantum computing include a quadratic improvement in searching unstructured databases Grover_1996 and an exponential enhancement in integer factorisation Shor_1994 ; Shor_1997 . Beyond these examples, quantum computing is expected to provide advantages across various domains, such as simulating chemical processes Cao2019 , solving optimisation problems jaderberg2022solving ; Symons_2023 , and advancing quantum machine learning phillipson2020quantum ; schuld2021machine .
@@
Quantum computing has the potential to also revolutionise the field of computational logistics and supply chain management by offering advantages in solving complex optimisation problems. Some of these problems include solving combinatorial optimisation problems, which are prevalent in logistics as we saw before: vehicle routing, facility location, and network design problems, which involve finding the best routes, locations, and structures to minimise costs and improve efficiency, and scheduling problems, such as production scheduling, job sequencing, and workforce management.
@@
Also the Quantum Technology and Application Consortium (QUTAC) bayerstadler2021industry acknowledges that optimisation and simulation problems are prevalent in the domain of production and logistics across industries, such as manufacturing, chemical and pharmaceutical production, insurance, and technology.. They state that real-world problems often involve numerous variables and constraints that classical algorithms struggle to address effectively, and that quantum optimisation approaches, such as quantum annealing and hybrid algorithms like the Quantum Approximate Optimisation Algorithm, are promising for their potential to provide higher quality solutions and faster solution times. Also Yarkoni et al. yarkoni2022quantum recognise scheduling and logistics as an important application area of quantum computing in their overview of industry applications.
@@
By reducing computation time significantly, quantum computing can help model and optimise the supply chain in real time, making it (more) resilient to disruptions, by finding alternative routes and sources during unexpected events, thus minimising the impact of disruptions. Quantum computers are also expected to be able to efficiently process and analyse vast amounts of data in the future, which is essential for demand forecasting, inventory management, and making real-time decisions in the supply chain. Lastly, quantum algorithms can help tackle multi-objective optimisation problems by simultaneously considering multiple conflicting objectives. This is beneficial for balancing cost reduction, service level improvement, and environmental sustainability. This is a great benefit for the two issues mentioned before.
@@
While quantum computing offers great promise for addressing logistics and supply chain optimisation problems, it is important to note that the technology is still in its early stages, and practical, large-scale implementations are limited. However, quantum hardware is evolving rapidly, and researchers are working on developing both the software stack needed and quantum algorithms for specific problems. Also quantum computing based approaches specifically tailored to logistics and supply chain management are studied and developed. A combination of classical and quantum computing techniques may be the most practical approach for solving logistics and supply chain challenges based on the current state of quantum computing hardware. This can be seen reflected in literature. A lot is happening in different research groups around the world. To get grip on this, in this paper an overview is given of over 80 published papers on this topic, both in detail and in the form of a compact overview. Next, conclusions from this overview are drawn and blind spots are recognised for further research. We continue the paper in Section 2 with an introduction to quantum computing and optimisation. Next, in Section 3, we will dive into the various areas on quantitative optimisation within the logistic and supply chain world and give an overview of the quantum approaches for the problems in those areas. As indicated we will end with a concise summary and recommendations for further research.
@@ 2 Quantum computing and optimisation
We can currently distinguish two paradigms in quantum computing: digital or gate-based computers/computing (GBC) and analog quantum computing, of which quantum annealers (QA) are an important example. GBC is most similar in operation to the current generation of computers. They are capable of performing operations (gate operations, such as AND, OR) on specific qubits or on multiple qubits at the same time. This allows for actual programming, which is often visualised via circuit diagrams. The QA, on the other hand, are single-purpose machines. Quantum annealing started with the work of Kadowaki and Nishimori kadowaki1998quantum . QA can basically do one thing only: find a minimum value of a specific function. This function is encoded in the qubits, after which a quantum mechanical evolution will lead to a solution that minimises the energy. In this section we will present GBC and QA in more detail and explain how they can be used in quantitative optimisation. First, we introduce some fundamental concepts in quantum computing.
@@ 2.1 Fundamental Concepts
Superposition, entanglement, interference and tunnelling are fundamental concepts in quantum mechanics, that underpin the power of quantum computing. Superposition refers to the ability of a quantum system to exist in multiple states simultaneously. Quantum computing uses qubits, or, quantum bits. Unlike classical bits, which can only represent either 0 or 1, a qubit can be in a superposition of both states. Mathematically, a qubit can be represented as a linear combination of the basis states |$\ket{0}$ and |$\ket{1}$, often denoted as $\alpha\ket{0}+\beta\ket{1}$, where $\alpha$ and $\beta$ are complex numbers that describe the probability amplitudes of each state. This notation is further explained in Section 2.2. When a measurement is made, the qubit collapses into one of the basis states with a probability determined by the square of the amplitudes, so it should hold that $|\alpha|^{2}+|\beta|^{2}=1$. Quantum parallelism is related to superposition. Quantum computational advantage is achieved by leveraging superposition in order to perform parallel computations on multiple qubits.
@@
Quantum interference is the third fundamental phenomenon in quantum mechanics, where the amplitudes of different quantum states combine in such a way, that they can reinforce or cancel each other out. It occurs when two or more quantum states, such as wave functions or probability distributions, overlap and interact with each other. Quantum interference is best understood through the concept of superposition. When superposed states interfere, the resulting probability distribution is not simply the sum of the individual probabilities. Instead, it depends on the relative phase of the states.
@@ 2.2 Gate based quantum computing
The first paradigm is digital or gate-based quantum computing. GBC rely more directly on qubits and gate operations. As said before, unlike classical bits, which can only represent 0 or 1, qubits can exist in a superposition of states, simultaneously representing multiple values. Through the application of quantum logic gates, which are analogous to classical logic gates, quantum computations are executed. These gates manipulate the quantum states of qubits, enabling operations such as superposition, entanglement, and interference. By leveraging a sequence of carefully crafted gate operations, quantum algorithms can be executed, allowing for the solution of problems that are intractable for classical computers. The precise control and manipulation of qubits, as well as the mitigation of errors due to decoherence, pose significant challenges in the practical implementation of gate-based quantum computing systems. The qubit states can be represented by matrices (vectors):
|$\ket{\psi}=\ket{0}=\binom{1}{0}\textrm{ and }\ket{\psi}=\ket{1}=\binom{0}{1}.$
@@
This means that a one-qubit quantum gate can be depicted as a unitary operator acting on a single qubit, which is represented as a two-dimensional system. A quantum gate acting on $n$ qubits is represented by a $2^{n}\times 2^{n}$ matrix. The state vector |$\ket{\psi}$ belongs to a Hilbert space $\mathcal{H}$, called the state space with vectors of length 1, using complex numbers. This state is given in Dirac notation: a quantum object state is represented by |$\ket{\psi}$, the ket of quantum state $\psi$. The bra of the same state vector, represented by $\bra{\psi}$ | is the conjugate transpose of the ket. If the vector contains only real numbers, the conjugate transpose is the same as the regular transpose, for a vector with complex entries, the conjugate transpose replaces each complex entry with its complex conjugate111if you have a complex number $z=a+bi$, where $a$ is the real part and $b$ is the imaginary part, then the complex conjugate of $z$, denoted as $\bar{z}$, is defined as $\bar{z}=a-bi.$ and then transposes the resulting matrix.
To give an example we look at the following. Figure 2 shows a quantum circuit for encoding the data: it encodes two data points in two qubits.
[…]
Figure 2: Amplitude encoding: two-qubit circuit for data encoding of two normalised two-dimensional data points.
A quantum circuits must be read from left to right. Each line in represents a qubit, in this case we have two qubits. Each rectangle is a gate operation, operating on one or two qubits. First, a Hadamard gate is applied to the first qubit to bring that qubit in a superposition state. Then a controlled rotation is used to encode the two data points $x_{i}=(x_{i0},x_{i1})$ for $i=0,1$ in the amplitudes of the second qubit. The angles of the rotations $\theta_{i}$ are chosen such that $\alpha_{i}=x_{i0}$ and $\beta_{i}=x_{i1}$. The initial state is now rotated to the desired state what can be expressed by:
[…]
Note that the first qubit acts as a counter, while the two features of the data point are encoded in the second qubit. Another option is to encode a single data point with four features, using just two qubits. Quantum programming languages can be used to construct and apply these circuits to quantum hardware. Examples of such languages are PyQuil koch2019introduction , QCL omer2005classical and Q# tolba2013q . These quantum programming languages only focus on that specific part of the quantum software stack van2019vision ; piattini2020talavera , whilst tools for other layers are also in development.
@@ 2.3 The QUBO
where $\mathbf{x}$ is a vector with binary decision variables and $Q$ is a square matrix with constant values. Many (constrained) combinatorial optimisation problems can easily be described in a QUBO formulation glover2019quantum . Take for example the problem
$\min\mathbf{c}^{T}\mathbf{x}\textrm{, under the constraints }A\mathbf{x}=\mathbf{b},$
@@
$\min\mathbf{c}^{T}\mathbf{x}+\lambda(A\mathbf{x}-\mathbf{b})^{T}(A\mathbf{x}-\mathbf{b}).$
@@
$\operatorname{argmin}\mathbf{x}^{T}P\mathbf{x}+\lambda(A\mathbf{x}-\mathbf{b})^{T}(A\mathbf{x}-\mathbf{b})=\mathbf{x}^{T}P\mathbf{x}+\mathbf{x}^{T}R\mathbf{x}+d=\mathbf{x}^{T}Q\mathbf{x},$
where the matrix $R$ and the constant $\mathbf{d}$ are the result of the multiplication and constant $\mathbf{d}$ can be neglected in the optimisation. Matrix $Q$ now depends on the penalty $\lambda$. The QUBO can then be reduced in a pre-processing step glover2018logical to create a smaller problem.
@@ 2.4 Quantum Annealing
The QUBO is given as input to the QA, which is an important representative of the second paradigm, analog quantum computing. The premise here is to create an equal superposition over all possible states of a collection of qubits. Then a problem-specific magnetic field, based on the QUBO formulation, is turned on, causing the qubits to interact with each other. Now the qubits move to the lowest energy state, from which the optimal solution of the original problem can be derived. Quantum annealing has similarities with the well-known Simulated Annealing (SA). However, where SA can only make thermal jumps, QA also uses thermal fluctuations and quantum fluctuations (like tunnelling, entanglement, etc). The most advanced QA is the version of D-Wave Systems (from now called D-Wave). They claim that their devices are practical implementations of adiabatic evolution farhi2000quantum . The evolution of a quantum state on the quantum processor of D-Wave is described by a time-dependent Hamiltonian ($H(t)$), consisting of the original Hamiltonian ($H_{0}$), whose ground state is easy to create, the equal superposition, and the final Hamiltonian ($H_{1}$), whose ground state encodes the solution of the current problem, via the QUBO. One specific linear annealing schedule is expressed by:
$H(t)=(1-\frac{t}{T})H_{0}+\frac{t}{T}H_{1}.$
@@
This system is initialised to the ground state of the original Hamiltonian, i.e. $H(0)=H_{0}$. The adiabatic theorem states that if the system evolves according to the Schrödinger equation, and the minimum spectral aperture of $H(t)$ is not zero, $H(T)$, for $T$ large enough, will converge to the ground state of $H_{1}$. Although we will not go into the technical details here, it is good to know that it is usually not possible to estimate an adequate time T for which the evolution to the desired state is assured. For some classes of problems, the optimal annealing time has been determined experimentally albash2018demonstration . There is therefore no guarantee of optimality. Furthermore, it is not trivial to find a good value for the penalty $\lambda$ and the so-called chainstrength; the QA is very sensitive to these parameters. Chainstrength is a parameter that results in a penalty if multiple qubits representing one variable do not have the same value in the solution. In addition, D-Wave’s state-of-the-art QA now has 5640 qubits with a connectivity of up to 15. This can model a fully connected clique problem with $n=177$. For problems with a lower dependency between the variables, larger problems can be solved. Problems that cannot be placed on a chip of this size will have to be broken into pieces, at the expense of the quality of the solution. D-Wave does offer a standard function that takes care of this decomposition automatically as well as offering hybrid computing pipelines. Note that in the remainder of this paper we use the term Quantum Annealing (QA) also if we actually mean the implemented Adiabatic Quantum Computing approach by D-Wave.
2.5 QAOA

### v2

@@ Abstract
The paper provides an overview of quantum approaches to routing, logistic network design, fleet maintenance, cargo loading, prediction, and scheduling problems. Notably, most solutions in the literature are hybrid, combining quantum and classical computing. The conclusion highlights the early stage of quantum computing, emphasising its potential impact on logistics and supply chain optimisation. In the final overview, the literature is categorised, identifying the dominance of quantum annealing and highlighting the need for more research in prediction and machine learning. The consensus is that quantum computing has great potential but faces current hardware limitations, necessitating further advancements for practical implementation.
@@
Quantum Computing Logistics Supply Chain Management Optimisation
@@ 1 Introduction
Quantitative optimisation is integral to logistics and supply chain management, driving efficiency, cost reduction, and overall performance improvement. Various challenges underscore the need for this field. These kinds of challenges and the corresponding optimisation problems are often divided into multiple levels, such as operational, tactical and strategic [23, 41], see Fig.  1. On an operational level, there is efficient route planning, as seen in vehicle routing problems (VRP), minimising transportation costs while meeting demand and adhering to constraints. Problems like stowage and cargo loading also need to be solved on an operational level. Next, in inventory management, a balance is sought between over-stocking and under-stocking, employing models like Economic Order Quantity (EOQ) for optimal order quantities. Production schedule optimisation addresses job scheduling, employee and machine scheduling, and production planning, minimising lead times and costs. Lastly, accurate demand forecasting is needed to optimise inventory, production, and transportation planning. On a tactical level, we have problems like fleet deployment, timetabling, and hub site design. On a strategic level, decisions on facility location, fleet size and mix, and network design, impact logistics costs.
In addition to these pure optimisation problems, there are two issues we also have to address in quantitative optimisation in this field. First, supply chain objectives often conflict, requiring multi-objective optimisation for informed decisions. Supplier selection involves balancing factors like cost, quality, lead times, and reliability. Risk management minimises disruptions from natural disasters or political events, ensuring operational continuity. Second, in logistics, we have to cope with the dynamic behaviour of the system. Dynamic optimisation adapts to real-time changes, crucial in a dynamic supply chain environment. Green supply chain optimisation focuses on sustainability, reducing environmental impact through carbon emission and energy usage minimisation.
@@
These issues make decision-making processes even more complex, considering various variables and constraints, ensuring a smooth flow of goods while maximising operational efficiency and minimising costs. Advanced techniques, including mathematical modelling and metaheuristics, are commonly used to tackle these challenges, which are computationally hard in most cases.
@@
Currently, there is a burgeoning computing paradigm known as quantum computing, which holds the potential to transform the contemporary computing landscape. Quantum computers, the devices central to this paradigm, leverage principles from quantum mechanics to achieve computational speeds and efficiencies that surpass the inherent capabilities of current classical computers for specific problems. The notable advancements attributable to quantum computing include a quadratic improvement in searching unstructured databases [58] and an exponential enhancement in integer factorisation [142, 143]. Beyond these examples, quantum computing is expected to provide advantages across various domains, such as simulating chemical processes [26], solving (or approximating) optimisation problems [68, 149], and advancing quantum machine learning [112, 136].
@@
Quantum computing also has the potential to revolutionise the field of computational logistics and supply chain management by offering advantages in solving complex optimisation problems better of faster. Some of these problems include solving combinatorial optimisation problems, which are prevalent in logistics as we saw before: vehicle routing, facility location, and network design problems, which involve finding the best routes, locations, and structures to minimise costs and improve efficiency, and scheduling problems, such as production scheduling, job sequencing, and workforce management. Note however that their ability to really solve NP-hard problems efficiently remains uncertain. Many combinatorial optimization problems, such as the Travelling Salesman Problem or Integer Programming, are NP-hard, meaning that no known polynomial-time algorithm exists for them on classical computers. Quantum computing operates within the complexity class BQP (Bounded-Error Quantum Polynomial Time), which includes problems that quantum computers can solve efficiently with high probability. However, NP-hard problems are generally believed to be outside BQP [39], suggesting that quantum computers may not provide exponential speedups for solving them exactly. However, quantum (assisted) heuristics may offer advantages in finding approximate solutions faster than classical methods.
@@
The Quantum Technology and Application Consortium (QUTAC) [14] also acknowledges that optimisation and simulation problems are prevalent in the domain of production and logistics across industries, such as manufacturing, chemical and pharmaceutical production, insurance, and technology. They state that real-world problems often involve numerous variables and constraints that classical algorithms struggle to address effectively, and that quantum optimisation approaches, such as quantum annealing (QA) and hybrid algorithms like the Quantum Approximate Optimisation Algorithm (QAOA), are promising for their potential to provide higher-quality solutions and faster solution times. Yarkoni et al. [169] also recognise scheduling and logistics as an important application area of quantum computing in their overview of industry applications.
@@
By reducing computation time significantly, quantum computing can help model and optimise the supply chain in real time, making it (more) resilient to disruptions, by finding alternative routes and sources during unexpected events, thus minimising the impact of disruptions. Quantum computers are also expected to efficiently process and analyse vast amounts of data in the future, which is essential for demand forecasting, inventory management, and making real-time decisions in the supply chain. Lastly, quantum algorithms can tackle multi-objective optimisation problems by simultaneously considering multiple conflicting objectives. This is beneficial for balancing cost reduction, service level improvement, and environmental sustainability, which is a great benefit for the two issues mentioned before.
@@
While quantum computing offers great promise for addressing logistics and supply chain optimisation problems, it is important to note that the technology is still in its early stages, and practical, large-scale implementations are limited. However, quantum hardware is evolving rapidly, and researchers are working on developing both the software stack needed and quantum algorithms for specific problems. Quantum computing-based approaches specifically tailored to logistics and supply chain management are also being studied and developed. A combination of classical and quantum computing techniques may be the most practical approach for solving logistics and supply chain challenges based on the current state of quantum computing hardware. This is reflected in literature as this work shows.
To get grip on this, this paper provides an overview of over 80 published papers on this topic, both in detail and in the form of a compact overview. Next, conclusions from this overview are drawn, and blind spots are recognised for further research. We continue the paper in Section 2 with an introduction to quantum computing and optimisation. Next, in Section 3, we will dive into the various areas of quantitative optimisation within the logistics and supply chain world and provide an overview of the quantum approaches for the problems in those areas. As indicated, we will end with a concise summary and recommendations for further research.
@@ 2 Quantum computing and optimisation
We can currently distinguish two paradigms in quantum computing: digital or gate-based computers/computing (GBC) and analogue quantum computing, of which quantum annealers are an important example. GBC is most similar in operation to the current generation of computers. They are capable of performing operations (gate operations, such as AND, OR) on specific qubits or on multiple qubits simultaneously. This allows for actual programming, which is often visualised through circuit diagrams. The QAs, on the other hand, are single-purpose machines. QA started with the work of Kadowaki and Nishimori [72]. QA can basically do one thing only: find the minimum value of a specific function. This function is encoded in the qubits, after which a quantum mechanical evolution leads to a solution that minimises the energy. In this section, we present GBC and QA in more detail and explain how they can be used in quantitative optimisation. First, we introduce some fundamental concepts in quantum computing.
@@ 2.1 Fundamental Concepts
Superposition, entanglement, interference, and tunnelling are fundamental concepts in quantum mechanics that underpin the power of quantum computing. Superposition refers to the ability of a quantum system to exist in multiple states simultaneously. Quantum computing uses qubits, or quantum bits. Unlike classical bits, which can only represent either 0 or 1, a qubit can be in a superposition of both states simultaneously. Mathematically, a qubit can be represented as a linear combination of the basis states |$\ket{0}$ and |$\ket{1}$, often denoted as $\alpha\ket{0}+\beta\ket{1}$, where $\alpha$ and $\beta$ are complex numbers that describe the probability amplitudes of each state. This notation is further explained in Section 2.2. When a measurement is made, the qubit collapses into one of the basis states with a probability determined by the square of the amplitudes, so it should hold that $|\alpha|^{2}+|\beta|^{2}=1$. Quantum parallelism is related to superposition. Quantum computational advantage is achieved by leveraging superposition to perform parallel computations on multiple qubits.
@@
Quantum interference is the third fundamental phenomenon in quantum mechanics, where the amplitudes of different quantum states combine in such a way that they can reinforce or cancel each other out. It occurs when two or more quantum states, such as wave functions or probability distributions, overlap and interact with each other. Quantum interference is best understood through the concept of superposition. When superposed states interfere, the resulting probability distribution is not simply the sum of the individual probabilities. Instead, it depends on the relative phase of the states.
@@ 2.2 Gate based quantum computing
The first paradigm is digital or gate-based quantum computing. GBC relies more directly on qubits and gate operations. As said before, unlike classical bits, which can only represent 0 or 1, qubits can exist in a superposition of states, simultaneously representing multiple values. Through the application of quantum logic gates, which are analogous to classical logic gates, quantum computations are executed. These gates manipulate the quantum states of qubits, enabling operations such as superposition, entanglement, and interference. By leveraging a sequence of carefully crafted gate operations, quantum algorithms can be executed, allowing for the solution of problems that are intractable for classical computers. The precise control and manipulation of qubits, as well as the mitigation of errors due to decoherence, pose significant challenges in the practical implementation of gate-based quantum computing systems. The qubit states can be represented by matrices (vectors):
|$\ket{\psi}=\ket{0}=\binom{1}{0}\textrm{ and }\ket{\psi}=\ket{1}=\binom{0}{1}.$ (1)
@@
This means that a one-qubit quantum gate can be depicted as a unitary operator acting on a single qubit, which is represented as a two-dimensional system. A quantum gate acting on $n$ qubits is represented by a $2^{n}\times 2^{n}$ matrix. The state vector |$\ket{\psi}$ belongs to a Hilbert space $\mathcal{H}$, called the state space with vectors of length 1, using complex numbers. This state is given in Dirac notation: a quantum object state is represented by |$\ket{\psi}$, the ket of quantum state $\psi$. The bra of the same state vector, represented by $\bra{\psi}$ | is the conjugate transpose of the ket. If the vector contains only real numbers, the conjugate transpose is the same as the regular transpose, for a vector with complex entries, the conjugate transpose replaces each complex entry with its complex conjugate111if you have a complex number $z=a+bi$, where $a$ is the real part and $b$ is the imaginary part, then the complex conjugate of $z$, denoted as $\bar{z}$, is defined as $\bar{z}=a-bi.$ and then transposes the resulting matrix. Quantum programming languages can be used to construct and apply quantum circuits to quantum hardware. Examples of such languages are PyQuil [77], QCL [105] and Q# [152]. These quantum programming languages only focus on that specific part of the quantum software stack [22, 115], whilst tools for other layers are also in development.
@@ 2.3 The QUBO
where $\mathbf{x}$ is a vector with binary decision variables and $Q$ is a square matrix with constant values. Many (constrained) combinatorial optimisation problems can easily be described in a QUBO formulation [55]. Take for example the problem
$\min\mathbf{c}^{T}\mathbf{x}\textrm{, under the constraints }A\mathbf{x}=\mathbf{b},$ (3)
@@
$\min\mathbf{c}^{T}\mathbf{x}+\lambda(A\mathbf{x}-\mathbf{b})^{T}(A\mathbf{x}-\mathbf{b}).$ (4)
@@
$\operatorname{argmin}\mathbf{x}^{T}P\mathbf{x}+\lambda(A\mathbf{x}-\mathbf{b})^{T}(A\mathbf{x}-\mathbf{b})=\mathbf{x}^{T}P\mathbf{x}+\mathbf{x}^{T}R\mathbf{x}+d=\mathbf{x}^{T}Q\mathbf{x},$ (5)
where the matrix $R$ and the constant $\mathbf{d}$ are the result of the multiplication and constant $\mathbf{d}$ can be neglected in the optimisation. Matrix $Q$ now depends on the penalty $\lambda$. The QUBO can then be reduced in a pre-processing step [56] to create a smaller problem.
@@ 2.4 Quantum Annealing
The QUBO is given as input to the QA, which is an important representative of the second paradigm, analogue quantum computing. The premise here is to create an equal superposition over all possible states of a collection of qubits. Then a problem-specific magnetic field, based on the QUBO formulation, is turned on, causing the qubits to interact with each other. Now the qubits move to the lowest energy state, from which the optimal solution of the original problem can be derived. QA has similarities with the well-known Simulated Annealing (SA). However, where SA can only make thermal jumps, QA also uses thermal fluctuations and quantum fluctuations (like tunnelling, entanglement, etc). The most advanced QA is the version of D-Wave Systems (from now called D-Wave). They claim that their devices are practical implementations of adiabatic evolution [48]. The evolution of a quantum state on the quantum processor of D-Wave is described by a time-dependent Hamiltonian ($H(t)$), consisting of the original Hamiltonian ($H_{0}$), whose ground state is easy to create, the equal superposition, and the final Hamiltonian ($H_{1}$), whose ground state encodes the solution of the current problem, via the QUBO. One specific linear annealing schedule is expressed by:
$H(t)=(1-\frac{t}{T})H_{0}+\frac{t}{T}H_{1}.$ (6)
@@
This system is initialised to the ground state of the original Hamiltonian, i.e. $H(0)=H_{0}$. The adiabatic theorem states that if the system evolves according to the Schrödinger equation, and the minimum spectral aperture of $H(t)$ is not zero, $H(T)$, for $T$ large enough, will converge to the ground state of $H_{1}$. Although we will not go into the technical details here, it is good to know that it is usually not possible to estimate an adequate time T for which the evolution to the desired state is assured. For some classes of problems, the optimal annealing time has been determined experimentally [5]. There is therefore no guarantee of optimality. Furthermore, it is not trivial to find a good value for the penalty $\lambda$ and the so-called chain strength; the QA is very sensitive to these parameters. Chain strength is a parameter that results in a penalty if multiple qubits representing one variable do not have the same value in the solution. In addition, D-Wave’s state-of-the-art annealer now has 5640 qubits with a connectivity of up to 15. This can model a fully connected clique problem with $n=177$ [97]. For problems with a lower dependency between the variables, larger problems can be solved. Problems that cannot be placed on a chip of this size will have to be broken into pieces, at the expense of the quality of the solution. D-Wave does offer a standard function that takes care of this decomposition automatically as well as offering hybrid computing pipelines. Note that in the remainder of this paper we use the term Quantum Annealing also if we actually mean the implemented Adiabatic Quantum Computing approach by D-Wave.
2.5 Quantum Approximate Optimisation Algorithm